OpenStudy (anonymous):

Can anyone explain to me what system of linear equations is? I will give medal and fan!!!!

OpenStudy (anonymous):

Hey you think you could help me???

OpenStudy (anonymous):

An equation that makes a straight line when it is graphed. Often written in the form: y = mx+b

OpenStudy (anonymous):

Yeah I kinda get that I just don't know how to solve its confusing can you help with a problem?

OpenStudy (anonymous):

yeah whats the problem

OpenStudy (anonymous):

Let me get it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Graph the system of inequalities . then use your graph to identify the point that represents a solution to the system. x>-2 y is less than or equal to 2x+7

OpenStudy (anonymous):

A. (-3,-1) B. (-1,4) C. (-1,6) D. (1,11)

OpenStudy (anonymous):

Do you get it? @samanthaself

OpenStudy (anonymous):

when you solve two equations at once for example y = 3x - 2 y = -6 - x

OpenStudy (anonymous):

a linear equation is in form y = mx + b , so a SYSTEM is more that 1 equation. usually two equations that you will solve at once using either the elimination or substitution method (:

OpenStudy (anonymous):

So how do I solve this one???

OpenStudy (anonymous):

@wendylisette can you help me?????

OpenStudy (anonymous):

are you sure its \[x > -2\] & \[y \le 2x + 7 ?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Why? @wendylisette

OpenStudy (anonymous):

well your answer will be between b c & d because x is greater than -2

OpenStudy (anonymous):

now plug in the x & y coordinates into the equation and see which one gives you the answer ,

OpenStudy (anonymous):

Is it C.? @wendylisette

OpenStudy (anonymous):

yes (: did you plug them in ?

OpenStudy (anonymous):

Yeah idk if I did it correctly but I got C. can you help me on another?

OpenStudy (anonymous):

@wendylisette

OpenStudy (anonymous):

sure , im on lunch break. ill try to help you before i leave

OpenStudy (anonymous):

Ok thanks!!!! I'll tag you in the question.

OpenStudy (anonymous):

I put it as another question and tagged you in it. :)

OpenStudy (anonymous):

@wendylisette